# Investigating Length Issues in Document-level Machine Translation

### v1

@@ A.1 The unifPE Algorithm
The unifPE algorithm sketched in Section 4.2 is formalised in Algorithm 1.
@@ A.3 Data Statistics and Other Details
The full data statistics are in Tables 7 and 8. All the full TED talks in our corpora start with the title, then the description and the talk before being split into pseudo-documents. Tags <description> and <title> are removed. When preparing our training and validation sets TED-U and TED-G (see Section 5.1), if concatenating the last sentence pair $(x_{n},y_{n})$ into the current pseudo-document pair exceeds $l_{max}$, $(x_{n},y_{n})$ will yield a single parallel sequence, to respect the maximum length $l_{max}$. The length distribution is illustrated in Figure 3.
@@
Table 7: Statistics of TEDTalks. All lengths are in English/French NLLB pieces.
@@
Table 8: Statistics of the test sets based on talks from IWSLT tst2014, tst2015, tst2016 and tst2017. “count” refers to the number of parallel pseudo-documents, “mean” denotes the average length of the source (i.e. English, top) or the reference (i.e. French, bottom) pseudo-documents. All lengths are counted in NLLB tokens.
256 512 768 1024 1200 1600 2048
NLLB 0.04 0.35 0.49 0.66 0.64 0.74 0.81
FT-U 0.01 0.03 0.08 0.09 0.13 0.26 0.44
Unif-U 0.01 0.03 0.07 0.10 0.20 0.34 0.32
FT-G 0.01 0.03 0.11 0.31 0.40 0.57 0.69
Unif-G 0.01 0.04 0.16 0.20 0.27 0.25 0.36
256 512 768 1024 1200 1600 2048
Tower 0.01 0.05 0.13 0.30 0.29 0.45 0.64
FT-U 0.01 0.05 0.10 0.15 0.13 0.23 0.59
Unif-U 0.02 0.05 0.09 0.15 0.16 0.28 0.61
FT-G 0.02 0.05 0.10 0.15 0.19 0.26 0.62
Unif-G 0.02 0.07 0.10 0.16 0.20 0.28 0.64
@@ A.4 Experimental Settings
For NLLB, we fine-tuned the pretrained model with learning rate $45e-$ 4, $500$ warm-up steps, $4$ parallel pseudo-documents per batch and $32$ gradient accumulation steps. An early stopping criterion with a patience of $5$ epochs is also applied, according to the d-BLEU evaluated on the validation set. For inference on test sets, the beam size is set to $5$ and the batch size is set to $4$.
@@ A.5 Detailed Evaluation Results
Table 11: Average difference (and p-values) in ds-BLEU or 100×100$\times$100 ×COMET between fine-tuned models (FT) and the corresponding pretrained models NLLB (top) and TowerBase (bottom). U and G denote the corpora TED-U and TED-G respectively. Positive values indicate that the fine-tuned model improves over the baseline.
@@
sent FT. 43.8 (0.98) 44.0 (0.99) 40.4 (1.00) 41.3 (1.00)
Unif. 42.4 (0.98) 42.5 (0.99) 39.7 (1.00) 40.2 (1.00)
256 FT. 43.5 (0.98) 43.1 (0.99) 40.4 (1.00) 40.8 (1.00)
Unif. 43.9 (0.98) 43.2 (0.99) 39.7 (1.00) 40.6 (1.00)
512 FT. 43.5 (0.98) 43.4 (0.98) 40.2 (1.00) 40.4 (1.00)
Unif. 43.8 (0.98) 43.8 (0.99) 39.8 (1.00) 41.0 (1.00)
768 FT. 36.6 (0.87) 36.4 (0.88) 35.3 (0.92) 35.6 (0.93)
Unif. 41.8 (0.95) 39.1 (0.88) 38.1 (0.95) 39.4 (0.97)
1024 FT. 28.6 (0.70) 29.1 (0.75) 28.9 (0.80) 28.7 (0.79)
Unif. 36.1 (0.81) 38.7 (0.87) 34.9 (0.87) 37.2 (0.92)
1200 FT. 25.2 (0.64) 25.8 (0.74) 24.1 (0.71) 24.3 (0.73)
Unif. 34.6 (0.80) 36.4 (0.82) 30.0 (0.74) 32.4 (0.80)
1600 FT. 18.2 (0.53) 19.3 (0.62) 19.2 (0.62) 19.6 (0.59)
Unif. 25.5 (0.59) 30.1 (0.70) 26.9 (0.68) 29.4 (0.72)
2048 FT. 15.4 (0.49) 12.4 (0.52) 16.7 (0.58) 14.8 (0.61)
Unif. 22.0 (0.51) 21.6 (0.50) 24.3 (0.60) 20.6 (0.53)
@@
sent FT. 84 85 85 84
Unif. 82 84 84 83
256 FT. 81 82 82 81
Unif. 81 82 82 81
512 FT. 81 82 81 80
Unif. 81 82 81 81
768 FT. 69 69 70 69
Unif. 74 75 74 73
1024 FT. 58 59 60 58
Unif. 67 65 67 67
1200 FT. 56 56 55 53
Unif. 61 65 62 60
1600 FT. 50 49 49 49
Unif. 55 58 54 56
2048 FT. 47 42 48 45
Unif. 51 49 51 48
Table 12: ds-BLEU (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-G (FT. ) and Unif-NLLB-G (Unif. ) trained on TED-G with target max source document length $M=2048$.
@@
sent FT. 44.2 (0.99) 43.5 (0.99) 40.4 (1.00) 41.4 (1.00)
Unif. 40.1 (0.95) 40.4 (0.97) 38.0 (1.00) 38.1 (0.99)
256 FT. 43.2 (0.98) 42.8 (0.99) 39.7 (1.00) 40.5 (1.00)
Unif. 42.8 (0.98) 42.4 (0.99) 39.5 (1.00) 40.3 (1.00)
512 FT. 42.9 (0.98) 43.1 (0.99) 39.2 (1.00) 39.4 (1.00)
Unif. 43.4 (0.98) 43.0 (0.99) 39.8 (1.00) 40.5 (1.00)
768 FT. 43.5 (0.98) 43.0 (0.99) 39.4 (1.00) 39.8 (1.00)
Unif. 44.0 (0.98) 43.7 (0.99) 40.4 (1.00) 40.9 (1.00)
1024 FT. 42.6 (0.96) 42.6 (0.97) 39.2 (1.00) 39.6 (1.00)
Unif. 42.6 (0.96) 44.1 (0.98) 39.6 (1.00) 40.6 (0.99)
1200 FT. 38.3 (0.88) 39.3 (0.91) 36.3 (0.92) 36.4 (0.92)
Unif. 39.5 (0.89) 43.0 (0.97) 38.0 (0.95) 39.2 (0.98)
1600 FT. 31.5 (0.77) 30.4 (0.73) 30.9 (0.83) 30.3 (0.80)
Unif. 31.5 (0.72) 34.4 (0.82) 34.2 (0.88) 33.8 (0.84)
2048 FT. 27.4 (0.69) 24.0 (0.63) 26.7 (0.75) 23.6 (0.71)
Unif. 30.2 (0.68) 25.3 (0.60) 31.5 (0.79) 28.0 (0.68)
@@
sent FT. 84 85 85 84
Unif. 82 83 83 83
256 FT. 80 81 81 81
Unif. 80 81 81 81
512 FT. 80 81 81 81
Unif. 81 82 81 81
768 FT. 80 81 81 81
Unif. 81 82 82 81
1024 FT. 78 79 78 78
Unif. 80 81 81 80
1200 FT. 71 73 70 69
Unif. 76 79 76 75
1600 FT. 61 60 61 60
Unif. 62 61 61 64
2048 FT. 57 53 57 54
Unif. 59 57 57 54
Table 13: ds-BLEU (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-U (FT. ) and Unif-NLLB-U (Unif. ) trained on TED-U with target max source document length $M=2048$.
@@
sent FT. 46.5 (0.98) 45.1 (0.99) 42.3 (1.00) 41.0 (1.00)
Unif. 46.5 (0.98) 45.0 (0.99) 42.3 (1.00) 41.1 (1.00)
256 FT. 44.6 (0.98) 45.1 (0.99) 42.3 (1.00) 41.9 (1.00)
Unif. 44.5 (0.99) 45.1 (0.99) 42.2 (1.00) 41.8 (1.00)
512 FT. 43.7 (0.98) 45.0 (1.00) 41.4 (1.00) 41.6 (1.00)
Unif. 43.7 (0.98) 44.8 (1.00) 41.4 (1.00) 41.4 (1.00)
768 FT. 44.0 (0.98) 44.2 (0.99) 40.3 (1.00) 40.7 (1.00)
Unif. 44.0 (0.98) 44.2 (0.99) 40.2 (1.00) 40.7 (1.00)
1024 FT. 42.7 (0.98) 40.6 (0.99) 38.9 (1.00) 40.4 (1.00)
Unif. 42.2 (0.97) 42.6 (0.99) 39.4 (1.00) 40.1 (1.00)
1200 FT. 42.6 (0.98) 43.0 (0.99) 38.9 (1.00) 40.8 (1.00)
Unif. 42.5 (0.98) 42.7 (0.99) 39.3 (1.00) 40.6 (1.00)
1600 FT. 42.0 (0.97) 41.1 (0.98) 37.0 (1.00) 38.7 (1.00)
Unif. 42.0 (0.97) 40.0 (0.98) 37.5 (1.00) 37.3 (1.00)
2048 FT. 33.0 (0.99) 32.5 (0.99) 31.6 (1.00) 29.2 (0.97)
Unif. 33.1 (0.99) 33.7 (0.99) 32.2 (0.98) 26.5 (0.97)
@@
sent FT. 85 86 85 85
Unif. 85 86 85 85
256 FT. 83 84 83 83
Unif. 83 84 83 83
512 FT. 82 84 83 82
Unif. 82 84 83 82
768 FT. 82 83 83 82
Unif. 82 83 83 82
1024 FT. 81 81 82 81
Unif. 81 82 83 81
1200 FT. 78 82 81 81
Unif. 78 82 81 81
1600 FT. 80 79 79 80
Unif. 79 78 79 78
2048 FT. 68 69 68 64
Unif. 69 70 70 62
Table 14: ds-BLEU (left) and 100×100$\times$100 ×COMET (right) scores for FT-Tower-G (FT. ) and Unif-Tower-G (Unif-full) trained on TED-G with target max source document length 2048 ($M=4096$).
@@
sent FT. 46.2 (0.99) 45.1 (0.99) 42.1 (1.00) 40.8 (1.00)
Unif. 46.4 (0.98) 45.1 (0.99) 42.2 (1.00) 40.9 (1.00)
256 FT. 46.3 (0.98) 45.1 (0.99) 42.3 (1.00) 41.8 (1.00)
Unif. 44.5 (0.98) 45.2 (0.99) 42.3 (1.00) 41.9 (1.00)
512 FT. 44.4 (0.98) 44.9 (0.99) 41.2 (1.00) 41.6 (1.00)
Unif. 43.0 (0.98) 45.0 (0.99) 41.4 (1.00) 41.6 (1.00)
768 FT. 44.1 (0.98) 44.6 (0.99) 41.1 (1.00) 40.8 (1.00)
Unif. 43.2 (0.99) 44.5 (0.99) 41.4 (1.00) 41.0 (1.00)
1024 FT. 43.9 (0.97) 41.6 (0.99) 39.6 (1.00) 39.7 (1.00)
Unif. 42.7 (0.98) 43.8 (0.99) 39.7 (1.00) 39.6 (1.00)
1200 FT. 43.2 (0.97) 42.9 (0.99) 39.5 (1.00) 40.8 (1.00)
Unif. 43.2 (0.98) 43.1 (0.99) 38.2 (1.00) 40.9 (1.00)
1600 FT. 41.6 (0.95) 40.5 (0.97) 37.6 (1.00) 39.8 (0.99)
Unif. 41.1 (0.97) 41.6 (0.98) 36.5 (1.00) 38.0 (1.00)
2048 FT. 34.4 (0.96) 35.3 (0.97) 31.2 (0.99) 28.2 (0.96)
Unif. 34.6 (0.97) 35.5 (0.98) 30.2 (1.00) 30.9 (0.97)

### v2

@@ A.1 The unifPE Algorithm
The unifPE algorithm briefly described in Section 4.2 is formalised in Algorithm 1.
@@ A.3 Data Statistics and Other Details
Full data statistics are given in Tables 8 and 9. All the full TED talks in our corpora start with the title, then the description and the talk before being split into pseudo-documents. <description> and <title> tags are removed. When preparing our training and validation sets TED-U and TED-G (see Section 5.1), if concatenating the last sentence pair $(x_{n},y_{n})$ into the current pseudo-document pair exceeds $l_{max}$, $(x_{n},y_{n})$ will yield a single parallel sequence, to respect the maximum length $l_{max}$. The length distribution is illustrated in Figure 3.
@@
Table 8: Statistics of the TED talks training and dev sets. ‘count’ denotes the number of parallel pseudo-documents. ‘mean’, ‘min’ and ‘max’ represent the average, minimum and maximum lengths of English/French pseudo-documents respectively, in NLLB tokens.
@@
Table 9: Statistics of the test sets based on talks from IWSLT tst2014, tst2015, tst2016 and tst2017 (see Section 5.1). ‘count’ refers to the number of parallel pseudo-documents. ‘mean’, ‘min’ and ‘max’ denote the average, minimum and maximum lengths of the source (i.e. English, top) or the reference (i.e. French, bottom) pseudo-documents. All lengths are counted in NLLB tokens.
@@ A.4 Experimental Settings
For NLLB, we fine-tuned the pretrained model with learning rate $5\mathrm{e}{-4}$, $500$ warm-up steps, $4$ parallel pseudo-documents per batch and $32$ gradient accumulation steps. An early stopping criterion with a patience of $5$ epochs is also applied, according to the d-BLEU evaluated on the validation set. For inference on test sets, the beam size is set to $5$ and the batch size is set to $4$.
@@ A.5 Detailed Evaluation Results
Table 11: Average difference (and p-values) in ds-BLEU or 100×100$\times$100 ×COMET between fine-tuned models (FT) and the corresponding pretrained models NLLB (top) and TowerBase (bottom). U and G denote the corpora TED-U and TED-G respectively. - for p-values $>0.05$. Positive values indicate that the fine-tuned model improves over the baseline.
@@
sent FT 43.8 (0.98) 44.0 (0.99) 40.4 (1.00) 41.3 (1.00)
Unif 42.4 (0.98) 42.5 (0.99) 39.7 (1.00) 40.2 (1.00)
SHAPE 40.7 (0.97) 41.5 (0.98) 38.3 (1.00) 39.4 (1.00)
256 FT 43.5 (0.98) 43.1 (0.99) 40.4 (1.00) 40.8 (1.00)
Unif 43.9 (0.98) 43.2 (0.99) 39.7 (1.00) 40.6 (1.00)
SHAPE 43.3 (0.98) 43.2 (0.99) 39.8 (1.00) 40.2 (0.99)
512 FT 43.5 (0.98) 43.4 (0.98) 40.2 (1.00) 40.4 (1.00)
Unif 43.8 (0.98) 43.8 (0.99) 39.8 (1.00) 41.0 (1.00)
SHAPE 40.6 (0.91) 40.5 (0.92) 37.0 (0.9) 40.4 (0.98)
768 FT 36.6 (0.87) 36.4 (0.88) 35.3 (0.92) 35.6 (0.93)
Unif 41.8 (0.95) 39.1 (0.88) 38.1 (0.95) 39.4 (0.97)
SHAPE 34.0 (0.75) 34.9 (0.77) 33.8 (0.84) 34.9 (0.85)
1024 FT 28.6 (0.70) 29.1 (0.75) 28.9 (0.80) 28.7 (0.79)
Unif 36.1 (0.81) 38.7 (0.87) 34.9 (0.87) 37.2 (0.92)
SHAPE 32.4 (0.73) 30.9 (0.69) 30.8 (0.75) 28.0 (0.69)
1200 FT 25.2 (0.64) 25.8 (0.74) 24.1 (0.71) 24.3 (0.73)
Unif 34.6 (0.80) 36.4 (0.82) 30.0 (0.74) 32.4 (0.80)
SHAPE 27.2 (0.61) 30.3 (0.68) 24.6 (0.64) 23.9 (0.60)
1600 FT 18.2 (0.53) 19.3 (0.62) 19.2 (0.62) 19.6 (0.59)
Unif 25.5 (0.59) 30.1 (0.70) 26.9 (0.68) 29.4 (0.72)
SHAPE 22.0 (0.50) 21.7 (0.50) 22.1 (0.56) 20.6 (0.57)
2048 FT 15.4 (0.49) 12.4 (0.52) 16.7 (0.58) 14.8 (0.61)
Unif 22.0 (0.51) 21.6 (0.50) 24.3 (0.60) 20.6 (0.53)
SHAPE 18.7 (0.43) 15.1 (0.44) 14.6 (0.38) 13.4 (0.48)
@@
sent FT 84 85 85 84
Unif 82 84 84 83
SHAPE 82 83 83 83
256 FT 81 82 82 81
Unif 81 82 82 81
SHAPE 80 82 81 80
512 FT 81 82 81 80
Unif 81 82 81 81
SHAPE 77 78 76 77
768 FT 69 69 70 69
Unif 74 75 74 73
SHAPE 68 69 68 68
1024 FT 58 59 60 58
Unif 67 65 67 67
SHAPE 60 62 61 59
1200 FT 56 56 55 53
Unif 61 65 62 60
SHAPE 55 59 55 54
1600 FT 50 49 49 49
Unif 55 58 54 56
SHAPE 51 52 50 48
2048 FT 47 42 48 45
Unif 51 49 51 48
SHAPE 46 43 47 44
Table 12: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-G (FT), Unif-NLLB-G (Unif), and SHAPE-NLLB-U (SHAPE) trained on TED-G with target max source document length $M=2048$.
@@
sent FT 44.2 (0.99) 43.5 (0.99) 40.4 (1.00) 41.4 (1.00)
Unif 40.1 (0.95) 40.4 (0.97) 38.0 (1.00) 38.1 (0.99)
SHAPE 39.4 (0.97) 40.0 (0.98) 36.3 (1.00) 37.8 (0.99)
256 FT 43.2 (0.98) 42.8 (0.99) 39.7 (1.00) 40.5 (1.00)
Unif 42.8 (0.98) 42.4 (0.99) 39.5 (1.00) 40.3 (1.00)
SHAPE 42.0 (0.97) 42.5 (0.99) 39.6 (1.00) 39.4 (1.00)
512 FT 42.9 (0.98) 43.1 (0.99) 39.2 (1.00) 39.4 (1.00)
Unif 43.4 (0.98) 43.0 (0.99) 39.8 (1.00) 40.5 (1.00)
SHAPE 39.7 (0.89) 41.1 (0.94) 38.2 (0.95) 39.0 (0.98)
768 FT 43.5 (0.98) 43.0 (0.99) 39.4 (1.00) 39.8 (1.00)
Unif 44.0 (0.98) 43.7 (0.99) 40.4 (1.00) 40.9 (1.00)
SHAPE 39.6 (0.88) 41.4 (0.93) 37.4 (0.93) 36.8 (0.91)
1024 FT 42.6 (0.96) 42.6 (0.97) 39.2 (1.00) 39.6 (1.00)
Unif 42.6 (0.96) 44.1 (0.98) 39.6 (1.00) 40.6 (0.99)
SHAPE 40.3 (0.91) 42.8 (0.96) 36.8 (0.92) 37.8 (0.94)
1200 FT 38.3 (0.88) 39.3 (0.91) 36.3 (0.92) 36.4 (0.92)
Unif 39.5 (0.89) 43.0 (0.97) 38.0 (0.95) 39.2 (0.98)
SHAPE 36.9 (0.84) 37.2 (0.87) 32.6 (0.82) 34.3 (0.86)
1600 FT 31.5 (0.77) 30.4 (0.73) 30.9 (0.83) 30.3 (0.80)
Unif 31.5 (0.72) 34.4 (0.82) 34.2 (0.88) 33.8 (0.84)
SHAPE 28.0 (0.68) 29.4 (0.68) 31.1 (0.79) 31.7 (0.82)
2048 FT 27.4 (0.69) 24.0 (0.63) 26.7 (0.75) 23.6 (0.71)
Unif 30.2 (0.68) 25.3 (0.60) 31.5 (0.79) 28.0 (0.68)
SHAPE 26.1 (0.63) 27.9 (0.66) 26.8 (0.69) 26.9 (0.74)
@@
sent FT 84 85 85 84
Unif 82 83 83 83
SHAPE 81 82 82 82
256 FT 80 81 81 81
Unif 80 81 81 81
SHAPE 79 81 81 80
512 FT 80 81 81 81
Unif 81 82 81 81
SHAPE 75 79 79 78
768 FT 80 81 81 81
Unif 81 82 82 81
SHAPE 75 78 75 75
1024 FT 78 79 78 78
Unif 80 81 81 80
SHAPE 74 78 75 73
1200 FT 71 73 70 69
Unif 76 79 76 75
SHAPE 68 70 69 66
1600 FT 61 60 61 60
Unif 62 61 61 64
SHAPE 57 59 62 59
2048 FT 57 53 57 54
Unif 59 57 57 54
SHAPE 51 52 57 53
Table 13: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-NLLB-U (FT) Unif-NLLB-U (Unif), and SHAPE-NLLB-U (SHAPE) trained on TED-U with target max source document length $M=2048$.
@@
sent FT 46.5 (0.98) 45.1 (0.99) 42.3 (1.00) 41.0 (1.00)
Unif 46.5 (0.98) 45.0 (0.99) 42.3 (1.00) 41.1 (1.00)
SHAPE 46.4 (0.98) 45.2 (0.99) 42.4 (1.00) 41.2 (1.00)
256 FT 44.6 (0.98) 45.1 (0.99) 42.3 (1.00) 41.9 (1.00)
Unif 44.5 (0.99) 45.1 (0.99) 42.2 (1.00) 41.8 (1.00)
SHAPE 46.2 (0.98) 45.2 (0.99) 42.4 (1.00) 42.1 (1.00)
512 FT 43.7 (0.98) 45.0 (1.00) 41.4 (1.00) 41.6 (1.00)
Unif 43.7 (0.98) 44.8 (1.00) 41.4 (1.00) 41.4 (1.00)
SHAPE 45.5 (0.98) 45.1 (0.99) 41.6 (1.00) 41.6 (1.00)
768 FT 44.0 (0.98) 44.2 (0.99) 40.3 (1.00) 40.7 (1.00)
Unif 44.0 (0.98) 44.2 (0.99) 40.2 (1.00) 40.7 (1.00)
SHAPE 45.6 (0.98) 44.4 (0.99) 41.3 (1.00) 41.5 (1.00)
1024 FT 42.7 (0.98) 40.6 (0.99) 38.9 (1.00) 40.4 (1.00)
Unif 42.2 (0.97) 42.6 (0.99) 39.4 (1.00) 40.1 (1.00)
SHAPE 44.5 (0.98) 40.9 (0.99) 39.2 (1.00) 39.7 (1.00)
1200 FT 42.6 (0.98) 43.0 (0.99) 38.9 (1.00) 40.8 (1.00)
Unif 42.5 (0.98) 42.7 (0.99) 39.3 (1.00) 40.6 (1.00)
SHAPE 44.0 (0.98) 42.8 (0.99) 39.3 (1.00) 40.6 (1.00)
1600 FT 42.0 (0.97) 41.1 (0.98) 37.0 (1.00) 38.7 (1.00)
Unif 42.0 (0.97) 40.0 (0.98) 37.5 (1.00) 37.3 (1.00)
SHAPE 42.2 (0.97) 40.6 (0.98) 38.5 (1.00) 39.2 (0.99)
2048 FT 33.0 (0.99) 32.5 (0.99) 31.6 (1.00) 29.2 (0.97)
Unif 33.1 (0.99) 33.7 (0.99) 32.2 (0.98) 26.5 (0.97)
SHAPE 34.1 (0.97) 35.1 (0.98) 32.1 (1.00) 30.2 (0.97)
@@
sent FT 85 86 85 85
Unif 85 86 85 85
SHAPE 85 86 85 85
256 FT 83 84 83 83
Unif 83 84 83 83
SHAPE 83 84 83 83
512 FT 82 84 83 82
Unif 82 84 83 82
SHAPE 82 84 83 82
768 FT 82 83 83 82
Unif 82 83 83 82
SHAPE 82 84 83 82
1024 FT 81 81 82 81
Unif 81 82 83 81
SHAPE 81 81 83 81
1200 FT 78 82 81 81
Unif 78 82 81 81
SHAPE 80 82 82 81
1600 FT 80 79 79 80
Unif 79 78 79 78
SHAPE 79 79 80 79
2048 FT 68 69 68 64
Unif 69 70 70 62
SHAPE 65 73 70 68
Table 14: ds-BLEU (and brevity penalty) (left) and 100×100$\times$100 ×COMET (right) scores for FT-Tower-G (FT), Unif-Tower-G (Unif) and SHAPE-Tower-G (SHAPE) trained on TED-G with target max source document length 2048 ($M=4096$).
@@
sent FT 46.2 (0.99) 45.1 (0.99) 42.1 (1.00) 40.8 (1.00)
Unif 46.4 (0.98) 45.1 (0.99) 42.2 (1.00) 40.9 (1.00)
SHAPE 46.3 (0.98) 45.2 (0.99) 42.4 (1.00) 41.0 (1.00)
256 FT 46.3 (0.98) 45.1 (0.99) 42.3 (1.00) 41.8 (1.00)
Unif 44.5 (0.98) 45.2 (0.99) 42.3 (1.00) 41.9 (1.00)
SHAPE 46.0 (0.98) 45.3 (0.99) 42.4 (1.00) 42.0 (1.00)
512 FT 44.4 (0.98) 44.9 (0.99) 41.2 (1.00) 41.6 (1.00)
Unif 43.0 (0.98) 45.0 (0.99) 41.4 (1.00) 41.6 (1.00)
SHAPE 44.5 (0.98) 45.0 (0.99) 41.3 (1.00) 41.6 (1.00)
768 FT 44.1 (0.98) 44.6 (0.99) 41.1 (1.00) 40.8 (1.00)
Unif 43.2 (0.99) 44.5 (0.99) 41.4 (1.00) 41.0 (1.00)
SHAPE 44.2 (0.99) 44.5 (0.99) 41.2 (1.00) 41.7 (1.00)
1024 FT 43.9 (0.97) 41.6 (0.99) 39.6 (1.00) 39.7 (1.00)
Unif 42.7 (0.98) 43.8 (0.99) 39.7 (1.00) 39.6 (1.00)
SHAPE 44.1 (0.98) 42.9 (0.99) 39.8 (1.00) 39.5 (1.00)
1200 FT 43.2 (0.97) 42.9 (0.99) 39.5 (1.00) 40.8 (1.00)
Unif 43.2 (0.98) 43.1 (0.99) 38.2 (1.00) 40.9 (1.00)
SHAPE 43.8 (0.98) 43.2 (0.99) 39.2 (1.00) 39.7 (1.00)
1600 FT 41.6 (0.95) 40.5 (0.97) 37.6 (1.00) 39.8 (0.99)
Unif 41.1 (0.97) 41.6 (0.98) 36.5 (1.00) 38.0 (1.00)
SHAPE 42.6 (0.96) 42.1 (0.98) 37.3 (1.00) 40.5 (1.00)
2048 FT 34.4 (0.96) 35.3 (0.97) 31.2 (0.99) 28.2 (0.96)
Unif 34.6 (0.97) 35.5 (0.98) 30.2 (1.00) 30.9 (0.97)
SHAPE 35.2 (0.97) 34.6 (0.95) 31.9 (0.99) 31.5 (0.96)